OpenStudy (anonymous):

A collection of dimes and nickels is worth $9.50. There are 132 coins in all. How many of each coin are there? ____dimes ____nickels @Directrix

Directrix (directrix):

Same song, fifth verse

OpenStudy (anonymous):

yea brb

Directrix (directrix):

Define variables. x will be the number of dimes y will be the number of nickels.

Directrix (directrix):

x + y = ?

OpenStudy (anonymous):

$9.50

Directrix (directrix):

No.

OpenStudy (anonymous):

x+y=132

Directrix (directrix):

ok

Directrix (directrix):

Each dime is worth 10 cents but the sum of the money is given in dollars. So, what is 10 cents written as part of a dollar. Hint, cent is a signal for hundredths.

Directrix (directrix):

Each dime is worth .10 x dimes are woth .10x You do the nickels in the same way.

OpenStudy (anonymous):

10x+5y=132

Directrix (directrix):

That is not what I asked.

Directrix (directrix):

Each nickel is worth ? y nickels are worth what times y

OpenStudy (anonymous):

5

Directrix (directrix):

No

OpenStudy (anonymous):

nickels=5 cents

Directrix (directrix):

5 cents = what part of a dollar

OpenStudy (anonymous):

5y

Directrix (directrix):

No

Directrix (directrix):

Each nickel is worth 5 cents which is $.05 y nickels is worth what?

OpenStudy (anonymous):

$9.50

Directrix (directrix):

NO

OpenStudy (anonymous):

sorry

Directrix (directrix):

.05*y is what the nickels are worth y nickels at five cents each.

OpenStudy (anonymous):

0.5y

OpenStudy (anonymous):

sorry meant .05y

Directrix (directrix):

x + y = 132, .10x + .05y = 9.50 Take these to the wolf.

OpenStudy (anonymous):

58 dimes 74 nickels

OpenStudy (anonymous):

thanks

Directrix (directrix):

Yes. The wolf threw up an approximately = symbol and I was seeing what the deal is on that.

Directrix (directrix):

Close, post, and move.

OpenStudy (anonymous):

@Directix I hate to be rude but you were a little rude